# Elysium: Exploring Object-level Perception in Videos via MLLM

### v1

@@ Abstract
Multi-modal Large Language Models (MLLMs) have demonstrated their ability to perceive objects in still images, but their application in video-related tasks, such as object tracking, remains understudied. This lack of exploration is primarily due to two key challenges. Firstly, extensive pretraining on large-scale video datasets is required to equip MLLMs with the capability to perceive objects across multiple frames and understand inter-frame relationships. Secondly, processing a large number of frames within the context window of Large Language Models (LLMs) can impose a significant computational burden. To address the first challenge, we introduce ElysiumTrack-1M, a large-scale video dataset paired with novel tasks: Referring Single Object Tracking (RSOT) and Video Referring Expression Generation (Video-REG). ElysiumTrack-1M contains 1.27 million annotated video frames with corresponding object boxes and descriptions. Leveraging this dataset, we conduct training of MLLMs and propose a token-compression model T-Selector to tackle the second challenge. Our proposed approach, Elysium: Exploring Object-level Perception in Videos via MLLM, is an end-to-end trainable MLLM that makes the first attempt to conduct object-level tasks in videos without requiring any additional plug-in or expert models. All codes and datasets are released at https://github.com/Hon-Wong/Elysium.
@@ 1 Introduction
In this paper, we introduce Elysium, an end-to-end trainable MLLM that is capable of handling both global-level and object-level tasks in the video field. To investigate the effectiveness of MLLMs on object-level tasks in videos, we introduce two novel tasks: Referring Single Object Tracking and Video Referring Expression Generation. Moreover, to address the issue of limited training data, we construct a large-scale dataset named ElysiumTrack-1M, which includes abundant trajectories and object descriptions to support tasks like Single Object Tracking (SOT), RSOT, and Video-REG. ElysiumTrack-1M is derived from the WebVid-10M dataset [5], and we employ a carefully designed processing pipeline to reduce noise in the generated labels of each video. To enable the MLLM to distinguish individual frames while reducing visual token use, we introduce a visual token compression network called T-Selector. This network offers a trade-off between performance and visual token consumption. We leave the exploration of additional object-level tasks, such as Multi-Object Tracking, Video Object Detection, Video Object Segmentation, and Referring Video Object Segmentation (RVOS) [29, 21, 57] as future work. Our contributions can be summarized in three key aspects:
1). We construct ElysiumTrack-1M, a million-scale dataset for object-level tasks in videos, supporting existing tasks like SOT and introducing new tasks like RSOT and Video-REG.
@@ 3 Construct ElysiumTrack-1M dataset
In this section, we will introduce two new tasks: RSOT and Video-REG. We will provide details on constructing the ElysiumTrack-1M dataset and discuss its scale and the tasks it can support. Additionally, we will address the evaluation of model performance and the metrics used on the ElysiumTrack-1M dataset.
@@ 3.1 Introduce RSOT and Video-REG Tasks
RSOT. Motivated by Referring Video Object Segmentation, we introduce a novel task called Referring Single Object Tracking. We define the task of RSOT as identifying and locating a specific object within an entire video by the given language expression. This task offers a more flexible format for tracking compared to traditional Single Object Tracking methods and establishes a valuable connection between language and tracking.
Video-REG. In addition to RSOT, we further extend the concept of Referring Expression Generation (REG) to the field of videos. We define this new task as Video-REG, which involves predicting a description of an object given its coordinates in any frame of a video. Unlike conventional REG tasks, Video-REG demands the model to possess temporal awareness. This is because the appearance of the object in the current frame might be affected by occlusion or motion blur, but it can be identified in other frames provided. Thus, the model needs to consider the temporal context to generate precise descriptions in Video-REG.
@@ 6 Conclusions and Limitations
Conclusions. In this paper, we present Elysium, an end-to-end trainable MLLM designed to leverage the full potential of MLLM in object-level perception tasks, encompassing both images and videos. We introduce two novel tasks, RSOT and Video-REG, which bridge the gap between language and tracking in videos. To facilitate these object-level tasks, we construct a large-scale dataset called ElysiumTrack-1M, providing support for tasks such as SOT, RSOT, and Video-REG. Furthermore, we propose a visual token compression network, T-Selector, to strike a balance between a large context window and overall performance. Through extensive experiments, we demonstrate that MLLM exhibits remarkable object perception abilities in videos. The results validate the effectiveness and potential of our proposed approach in leveraging MLLM for object-level perception tasks.

### v2

@@ Abstract
Multi-modal Large Language Models (MLLMs) have demonstrated their ability to perceive objects in still images, but their application in video-related tasks, such as object tracking, remains understudied. This lack of exploration is primarily due to two key challenges. Firstly, extensive pretraining on large-scale video datasets is required to equip MLLMs with the capability to perceive objects across multiple frames and understand inter-frame relationships. Secondly, processing a large number of frames within the context window of Large Language Models (LLMs) can impose a significant computational burden. To address the first challenge, we introduce ElysiumTrack-1M, a large-scale video dataset supported for three tasks: Single Object Tracking (SOT), Referring Single Object Tracking (RSOT), and Video Referring Expression Generation (Video-REG). ElysiumTrack-1M contains 1.27 million annotated video frames with corresponding object boxes and descriptions. Leveraging this dataset, we conduct training of MLLMs and propose a token-compression model T-Selector to tackle the second challenge. Our proposed approach, Elysium: Exploring Object-level Perception in Videos via MLLM, is an end-to-end trainable MLLM that attempts to conduct object-level tasks in videos without requiring any additional plug-in or expert models. All codes and datasets are released at https://github.com/Hon-Wong/Elysium.
@@ 1 Introduction
In this paper, we introduce Elysium, an end-to-end trainable MLLM that is capable of handling both global-level and object-level tasks in the video field. To address the issue of limited training data, we construct a large-scale dataset named ElysiumTrack-1M, which includes abundant trajectories and object descriptions to support tasks like Single Object Tracking (SOT), RSOT, and Video-REG. ElysiumTrack-1M is derived from the WebVid-10M dataset [5], and we employ a carefully designed processing pipeline to reduce noise in the generated labels of each video. To enable the MLLM to distinguish individual frames while reducing visual token use, we introduce a visual token compression network called T-Selector. This network offers a trade-off between performance and visual token consumption. We leave the exploration of additional object-level tasks, such as Multi-Object Tracking, Video Object Detection, Video Object Segmentation, and Referring Video Object Segmentation (RVOS) [29, 21, 57] as future work. Our contributions can be summarized in three key aspects:
1). We construct ElysiumTrack-1M, a million-scale dataset for object-level tasks in videos, supporting existing tasks like SOT and introducing benchmarks for RSOT and Video-REG.
@@ 3 Construct ElysiumTrack-1M dataset
In this section, we will give a clear definition about two tasks: RSOT and Video-REG. We will provide details on constructing the ElysiumTrack-1M dataset and discuss its scale and the tasks it can support. Additionally, we will address the evaluation of model performance and the metrics used on the ElysiumTrack-1M dataset.
@@ 3.1 Introduce RSOT and Video-REG Tasks
RSOT. Motivated by Referring Video Object Segmentation, we define the task of RSOT as identifying and locating a specific object within an entire video by the given language expression only (no positional cue is used). This task offers a more flexible format for tracking compared to traditional Single Object Tracking methods and establishes a valuable connection between language and tracking.
Video-REG. In addition to RSOT, we further extend the concept of Referring Expression Generation (REG) to the field of videos. We define this task as Video-REG, which involves predicting a description of an object given its coordinates in any frame of a video. Unlike conventional REG tasks, Video-REG demands the model to possess temporal awareness. This is because the appearance of the object in the current frame might be affected by occlusion or motion blur, but it can be identified in other frames provided. Thus, the model needs to consider the temporal context to generate precise descriptions in Video-REG.
@@ 6 Conclusions and Limitations
Conclusions. In this paper, we present Elysium, an end-to-end trainable MLLM designed to leverage the full potential of MLLM in object-level perception tasks, encompassing both images and videos, paired with two tasks, RSOT and Video-REG, which bridge the gap between language and tracking in videos. To facilitate these object-level tasks, we construct a large-scale dataset called ElysiumTrack-1M, providing support for tasks such as SOT, RSOT, and Video-REG. Furthermore, we propose a visual token compression network, T-Selector, to strike a balance between a large context window and overall performance. Through extensive experiments, we demonstrate that MLLM exhibits remarkable object perception abilities in videos. The results validate the effectiveness and potential of our proposed approach in leveraging MLLM for object-level perception tasks.